OpenStudy (anonymous):

one line passes through the points G(-1,5)and H (2,9). a second line passes through the points x(-3.-2) and y (5,12). which line has a greater slope? explain how you can tell by graphing the two lines. explain how you can tell bby calculating the slopes of the lines

OpenStudy (anonymous):

ok so by graphing you would need to graph the points properly and then draw a line going through the points GH and XY. From here you look at which line has a greater "slope" meaning that you look at the line that has a higher angle from either the x or -x axis. this will make more sense once we get into the actual solution. So the actual solution is the following: For finding the slope we do the following ... slope = rise / run = G(rise) - H(rise) / G(run) - H(run) = 5 - 9 / -1 -2 = -4 / -3 = 4/3. So you can plot the line y = (4/3)x on the graphing paper. Now to find the slope of the second line. slope = rise / run = x(rise) - y(rise) / x(run) - y(run) = -2 - 12 / -3 -5 = -14/-8 = 14/8 = 7/4. So you can plot the line y = (7/4)x on the graphing paper. Now if you actually put those two lines on the same graph you will see that the line with the slope of (7/4) is steeper than the line of the slope (4/3). Therefore, you can say that line y = (7/4)x has a greater slope than line y = (4/3)x